### v1

@@ Abstract
The celebrated conjecture by Payne, Pólya and Weinberger (1956) states that for the fixed membrane problem, the ratio of the first two eigenvalues, $\lambda_{2}/\lambda_{1}$, is maximized by a disk. A more general dimensional version of this conjecture was later resolved by Ashbaugh and Benguria in the 1990s. For the Robin Laplacian, Payne and Schaefer (2001) formulated an analogous conjecture, positing that the ratio $\mu_{2}/\mu_{1}$ is also maximized by a disk for a range of the boundary parameter $\sigma$. This was later restated by Henrot in 2003. In this work, we affirm this conjecture for all dimensions $N\geq 2$ and for all $\sigma$ greater than or equal to a critical positive value $\sigma_{*}$ which may depend on $\Omega$. Furthermore, we prove that the maximum value of $\mu_{2}/\mu_{1}$ is strictly decreasing in $\sigma$ over the entire interval $(0,+\infty)$. Our result provides a positive answer to a variant of Yau’s Problem 77: by measuring the ratio of the first two eigenfrequencies, one can determine whether an elastically supported drum is circular.
@@ 1 Introduction
The upper bound of $3$ for this ratio has been progressively improved by several mathematicians. Notably, Brands [11] improved it to $2.686$, de Vries [46] further advanced it to $2.658$, and Chiti [15] ultimately reached $2.586$. In higher dimensions $N\geq 2$, the generalized PPW conjecture states that
@@
This conjecture was conclusively proved in the early 1990s by Ashbaugh and Benguria [6, 7, 8]. For convenience, we will refer to this result as the Ashbaugh-Benguria inequality in the subsequent discussion
@@
For general dimensions $N\geq 2$, a natural extension of this problem is to determine whether the ratio $\mu_{2}/\mu_{1}$ is maximized by an $N$-dimensional ball for all $\sigma>0$, or at least for a certain range of $\sigma$. To investigate this, let $B_{R}$ denote the ball of radius $R\in(0,+\infty)$ centered at the origin in $\mathbb{R}^{N}$, and let $\nu=N/2-1$.
In this work, we establish the validity of this conjecture for all dimensions $N\geq 2$ and for all sufficiently large values of the boundary parameter $\sigma$.
@@
Theorem 1.1. Let $\Omega\subset\mathbb{R}^{N}$ ($N\geq 2$) be a given bounded domain whose boundary satisfies an interior sphere condition. There exists a constant $\sigma_{*}(\Omega)>0$ such that for all $\sigma\geq\sigma_{*}$, the ratio of the first two Robin eigenvalues of the Laplacian on $\Omega$ satisfies
$\frac{\mu_{2}}{\mu_{1}}\leq\frac{\mu_{2}}{\mu_{1}}\Big|_{B_{R}}=\frac{k_{\nu+1,1}^{2}}{k_{\nu,1}^{2}},$
@@
where $B_{R}$ is a ball of radius $R$ chosen to have the same volume as $\Omega$ and $k_{\nu+l,m}$ denotes the $m$th positive zero of the function $kJ_{\nu+l+1}(k)-(\sigma+l)J_{\nu+l}(k)$. Furthermore, equality holds if and only if $\Omega$ itself is a ball ($\Omega=B_{R}$).
@@
Theorem 1.1 has a direct physical implication for a drum with an elastically supported boundary: one can indeed determine whether such a drum is circular from its first two eigenfrequencies. Specifically, if the ratio $\mu_{2}/\mu_{1}$ attains the universal upper bound $k_{1,1}^{2}/k_{0,1}^{2}$ which is independent of the shape and size of the drum, then the drum must be circular. Conversely, any deviation of this ratio from the maximum value signifies a non-circular shape. Moreover, the ratio of the first two tones cannot exceed $k_{1,1}^{2}/k_{0,1}^{2}$. This conclusively affirms the claim in Yau’s Problem 77 [48] for the case of an elastic drum.
@@
Our result also provides a nuanced answer to Kac’s famous question [31], “Can one hear the shape of a drum?” While the general answer is negative, our theorem shows that for an elastically supported drum, one can definitively “hear” its circularity from the first two tones.
@@
Theorem 1.2. The maximum value of $\mu_{2}/\mu_{1}$ is strictly decreasing with respect to $\sigma\in(0,+\infty)$. Furthermore, the function $\sigma\mapsto k_{\nu+1,1}^{2}/k_{\nu,1}^{2}$ is continuously and strictly decreasing on $(0,+\infty)$, and its range spans from $\infty$ (as $\sigma\to 0^{+}$) down to $j_{\nu+1,1}^{2}/j_{\nu,1}^{2}$ (as $\sigma\to+\infty$).
@@
Figure 1 provides a schematic illustration of the ratio $k_{\nu+1,1}^{2}/k_{\nu,1}^{2}$ as a function of $\sigma$. A crucial implication of Theorem 1.2 for $N=2$ is the existence of a unique critical value $\sigma^{*}>0$ such that $k_{\nu+1,1}^{2}/k_{\nu,1}^{2}\left(\sigma^{*}\right)=3$. Furthermore, the ratio exceeds $3$ for $\sigma<\sigma^{*}$ and falls below $3$ for $\sigma>\sigma^{*}$. Consequently, when $\Omega$ is a ball, the inequality $\mu_{2}/\mu_{1}\leq 3$ holds if and only if $\sigma\geq\sigma^{*}$. This result precisely identifies the optimal parameter range, thereby refining the earlier estimate in [39].
@@
where $\gamma=\sqrt{\lambda_{1}}/\alpha$. Chiti [13, 14, 15] proved—for a class of equations more general than the Laplacian—that there exists $s_{1}\in\left(0,\left|B_{1/\gamma}\right|\right)$ such that the decreasing rearrangements $\widetilde{u}_{1}^{*}(s)$ and $\widetilde{z}^{*}(s)$ (defined in Section 3) satisfy
$\left\{\begin{array}[]{ll}\widetilde{z}^{*}(s)\geq\widetilde{u}_{1}^{*}(s)&\text{for}\,\,\,\left[0,s_{1}\right],\\ \widetilde{z}^{*}(s)\leq\widetilde{u}_{1}^{*}(s)&\text{for}\,\,\,\left[s_{1},\left|B_{1/\gamma}\right|\right],\end{array}\right$.
where $|B_{1/\gamma}|=\text{meas}\left(B_{1/\gamma}\right)$. This result was later extended by Ashbaugh and Benguria [7, Theorem A.2] to a broader class of second-order elliptic equations.
@@
$\left(u_{1}^{*}\right)^{\prime}(s)\leq\mu_{1}N^{-2}C_{N}^{-2/N}s^{\frac{2}{N}-2}\int_{0}^{s}u_{1}^{*}(t)\,\text{d}t,$
where $C_{N}=\pi^{N/2}/\Gamma\left(1+N/2\right)$ is the volume of the $N$-dimensional unit ball. The principal difficulty in deriving this inequality for the Robin problem—and indeed a central challenge in applying symmetric rearrangement methods to it—stems from the behavior of the eigenfunction on the boundary. Unlike the Dirichlet case, $u_{1}$ is generally non-constant on $\partial\Omega$. Consequently, the level set $\left\{x\in\Omega:u_{1}>t\right\}:=U_{t}$ may include portions of $\partial\Omega$ where $u_{1}>t$, which fundamentally alters its perimeter and geometric properties. This complication prevents a direct application of the techniques used for the fixed membrane problem [7, 13, 43], as noted in [2].
To overcome this challenge, we adapt an elegant idea from Alvino et al. [2, 3], which involves decomposing the boundary $\partial U_{t}$ of a level set into its internal part and its intersection with $\partial\Omega$. By selecting a sufficiently large boundary parameter $\sigma$, we can ensure that a specific threshold level exceeds the maximum value of $u_{1}$ on $\partial\Omega$. This allows us to exclude the external part of $\partial U_{t}$, thereby establishing the required isoperimetric-type inequality on a suitable subinterval—a result sufficient for our purposes. Building on this, we then derive a new Chiti-type comparison theorem. All these developments are presented in Section 3. Furthermore, leveraging the Faber-Krahn inequality for the Robin problem (established by Bossel [10] for $N=2$ and Daners [18] for $N\geq 2$), we also prove a key domain comparison result in this section, which serves as another essential prerequisite for the proof of Theorem 1.1.
The proof of Theorem 1.1 is presented in Section 4. Owing to the fundamental differences between our new Chiti-type comparison result and the classical version in [7, Theorem A.2], the structure of our proof diverges significantly from the original approach. In particular, establishing the necessity of the equality condition requires more complex and refined arguments. Finally, the proof of Theorem 1.2 is completed in the last section.
@@ 2 Strict monotonicity of $w(r)$ and $B(r)$
where the first equality follows from the Bessel function identity $J_{\nu+1}(x)+J_{\nu+3}(x)=2(\nu+2)J_{\nu+2}(x)/x$, and the final inequality holds because $\beta<j_{\nu+1,1}$ ensures the integrand remains positive. This completes the proof.∎
@@
We conclude this section by analyzing the asymptotic behavior of the first eigenfunction with respect to the boundary parameter $\sigma$. Let $u_{1}$ be any positive eigenfunction corresponding to $\mu_{1}$, and denote $u_{1,pM}=\max_{\partial\Omega}u_{1}$ and $u_{1,M}=\max_{\Omega}u_{1}$. Since $\partial\Omega$ satisfies an interior sphere condition, the Hopf Lemma [25, Lemma 3.4] implies that $u_{1,pM}<u_{1,M}$.
@@ 3 Chiti type comparison results
$\int_{\Omega}u^{2}(x)\,\text{d}x=\int_{0}^{M}\left(u^{*}\right)^{2}(s)\,\text{d}s=\int_{\Omega^{*}}\left(u^{\star}\right)^{2}(x)\,\text{d}x,$ (3.1)
@@
Let $R=\alpha/\sqrt{\mu_{1}}$. Since $\alpha^{2}$ is the first eigenvalue of the Robin problem on the unit ball $B$, a standard scaling argument shows that $\mu_{1}$ is the first eigenvalue of the corresponding problem on the ball $B_{R}$
@@
where $c$ is a nonzero positive constant. Given that $\alpha=k_{\nu,1}$, we have $z(R)>0$, which contrasts with the Dirichlet case where the eigenfunction vanishes on the boundary.
@@
By the Faber-Krahn inequality for the Robin Laplacian—established by Bossel [10] for $N=2$ and Daners [18] for $N\geq 2$—we have
@@
It follows that $\mu_{1}\left(B_{R}\right)=\mu_{1}\left(\Omega^{*}\right)$. So one has that $B_{R}=\Omega^{*}$ which is a contradiction. Hence, $B_{R}\subseteq\Omega^{*}$ or $\left|B_{R}\right|\leq\left|\Omega^{*}\right|=|\Omega|$. Furthermore, if $\left|B_{R}\right|=|\Omega|$, then $B_{R}=\Omega^{*}$, and thus
@@
By Lemma 3.2, $\left|B_{R}\right|=|\Omega|$ if and only if $\Omega=B_{R}$. In this case, the conclusion of Theorem 1.1 holds trivially for all $\sigma>0$. We may therefore assume throughout the remainder of the paper that $\left|B_{R}\right|<|\Omega|$.
@@
Lemma 3.3. For $\sigma\geq\sigma_{*}$ with some $\sigma_{*}>0$, if $z^{*}(0)<u_{1}^{*}(0)$, then $z^{*}(s)<u_{1}^{*}(s)$ for all $s\in\left[0,\left|B_{R}\right|\right)$.
@@
By Proposition 2.3 and the right-continuity of $\mu(t)$, we have
$\lim_{\sigma\rightarrow+\infty}\mu\left(u_{1,pM}\right)=\lim_{\sigma\rightarrow+\infty}\mu\left(u_{1,m}\right)=|\Omega|.$
Since $|\Omega|>\left|B_{R}\right|$, there exists $\sigma_{*}>0$ such that
@@
for all $\sigma\geq\sigma_{*}$. By Lemma 3.1 and the definition of $v$, we have
@@
for almost every $s\in\left[0,s_{1}\right]$. Combining this with inequality (3.9), we conclude that the same inequality holds for almost every $s\in\left[0,\left|B_{R}\right|\right]$. We need to point out that the constant $\sigma_{*}$ is not universal but depends on the domain [^2] It is our pleasure to thank Professor Richard Laugesen for bringing this to our attention..
@@
$\int_{0}^{M}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s.$
@@
$\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s<\int_{0}^{\left|B_{R}\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s\leq\int_{0}^{M}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s,$ (3.10)
@@
Lemma 3.4. When $\sigma\geq\sigma_{*}$ and $z^{*}(0)=u_{1}^{*}(0)$, the following holds: either $z^{*}(s)\geq u_{1}^{*}(s)$ for all $s\in\left[0,\left|B_{R}\right|\right]$, or there exists some $s_{1}\in\left(0,\left|B_{R}\right|\right)$ such that $z^{*}(s)\geq u_{1}^{*}(s)$ for $s\in\left[0,s_{1}\right]$ and $z^{*}(s)\leq u_{1}^{*}(s)$ for $s\in\left(s_{1},\left|B_{R}\right|\right]$.
@@
$\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s<\int_{0}^{\left|B_{R}\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s\leq\int_{0}^{M}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s,$
@@
Theorem 3.1. Let (3.5) hold. For $\sigma\geq\sigma_{*}$, the functions $u_{1}^{*}(s)$ and $z^{*}(s)$ satisfy either
@@
(I) if $f=f(|x|)$ is nonnegative and decreasing (or increasing) as a function of $r=|x|$ for $x\in\Omega$, then $f^{\star}(r)\leq f(r)$ (or $f_{\star}(r)\geq f(r)$) for $r$ between 0 and the radius of $\Omega^{*}$, with strict inequality for $r\in\left(\min_{x\not\in}|x|,r^{*}\right)$ if $f$ is strictly monotone there, where $r^{*}$ denote the radius of $\Omega^{*}$;
@@
$f^{\star}(r)=f(r)\,\,\,\text{for}\,\,\,r\in\left[0,r^{*}\right].$
@@
$f_{\star}(r)=f(r)\,\,\,\text{for}\,\,\,r\in\left[0,r^{*}\right].$
@@
By Theorem 3.1, when $\sigma\geq\sigma_{*}$, the functions $u_{1}^{\star}(r)$ and $z(r)$ satisfy either
@@
$\left.-\int_{R}^{r^{*}}f(r)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right).$
@@
$\leq\,$ $NC_{N}\left(\int_{0}^{R}f(R)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{r^{*}}f(R)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $NC_{N}f(R)\left(\int_{0}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$

### v2

@@ Abstract
The celebrated conjecture by Payne, Pólya and Weinberger (1956) states that for the fixed membrane problem, the ratio of the first two eigenvalues, $\lambda_{2}/\lambda_{1}$, is maximized by a disk. A more general dimensional version of this conjecture was later resolved by Ashbaugh and Benguria in the 1990s. For the Robin Laplacian, Payne and Schaefer (2001) formulated an analogous conjecture, positing that the ratio $\mu_{2}/\mu_{1}$ is also maximized by a disk for a range of the boundary parameter $\sigma$. This was later restated by Henrot in 2003. In this work, under some suitable conditions, we affirm this conjecture for all dimensions $N\geq 2$ and for all $\sigma>0$. Furthermore, we prove that the maximum value of $\mu_{2}/\mu_{1}$ is strictly decreasing in $\sigma$ over the entire interval $(0,+\infty)$. Our result provides a positive answer to a variant of Yau’s Problem 77: by measuring the ratio of the first two eigenfrequencies, one can determine whether an elastically supported drum is circular.
@@ 1 Introduction
The upper bound $3$ for this ratio has been progressively improved by several mathematicians. Notably, Brands [11] improved it to $2.686$, de Vries [46] further advanced it to $2.658$, and Chiti [15] ultimately reached $2.586$. In higher dimensions $N\geq 2$, the generalized PPW conjecture states that
@@
This conjecture was conclusively proved in the early 1990s by Ashbaugh and Benguria [6, 7, 8]. For convenience, we will refer to this result as the Ashbaugh-Benguria inequality in the subsequent discussion.
@@
For general dimensions $N\geq 2$, a natural extension of this problem is to determine whether the ratio $\mu_{2}/\mu_{1}$ is maximized by an $N$-dimensional ball for all $\sigma>0$, or at least for a certain range of $\sigma$. To investigate this conjecture, we first introduce some notations. Let $B_{R}$ denote the ball of radius $R\in(0,+\infty)$ centered at the origin in $\mathbb{R}^{N}$, and set $\nu=N/2-1$. Let $u_{1}$ be any given positive eigenfunction corresponding to $\mu_{1}$, and denote $u_{1,pM}=\max_{\partial\Omega}u_{1}$ and $u_{1,M}=\max_{\Omega}u_{1}$. If $\partial\Omega$ satisfies the interior sphere condition, the Hopf Lemma [25, Lemma 3.4] implies that $u_{1,pM}<u_{1,M}$. Let $\mu(t)$ denote the distribution function of $u_{1}$ (see the exact defition in Section 3). We also let $R=\alpha/\sqrt{\mu_{1}}$ and $\widetilde{R}=\left(\mu\left(u_{1,pM}\right)/C_{N}\right)^{1/N}$ where $C_{N}=\pi^{N/2}/\Gamma\left(1+N/2\right)$ is the volume of the $N$-dimensional unit ball.
The first main result is the following theorem.
@@
Theorem 1.1. Let $\Omega\subset\mathbb{R}^{N}$ ($N\geq 2$) be a given bounded domain whose boundary satisfies an interior sphere condition. If $\widetilde{R}\geq R$, the ratio of the first two Robin eigenvalues of the Laplacian on $\Omega$ satisfies
$\frac{\mu_{2}}{\mu_{1}}\leq\frac{\mu_{2}}{\mu_{1}}\Big|_{B_{R_{*}}}=\frac{k_{\nu+1,1}^{2}}{k_{\nu,1}^{2}}$
@@
for all $\sigma>0$, where $B_{R_{*}}$ is a ball of radius $R_{*}$ chosen to have the same volume as $\Omega$ and $k_{\nu+l,m}$ denotes the $m$th positive zero of the function $kJ_{\nu+l+1}(k)-(\sigma+l)J_{\nu+l}(k)$. Furthermore, equality holds if and only if $\Omega$ itself is a ball ($\Omega=B_{R_{*}}$).
@@
Under the assumptions of Theorem 1.1, our conclusion indicates that the above PPW type conjecture is true for all dimensions $N\geq 2$ and for all $\sigma>0$. In particular, Theorem 1.1 has a direct physical implication for a drum with an elastically supported boundary: one can indeed determine whether such a drum is circular from its first two eigenfrequencies. Specifically, under our conditions, if the ratio $\mu_{2}/\mu_{1}$ attains the universal upper bound $k_{1,1}^{2}/k_{0,1}^{2}$ which is independent of the shape and size of the drum, then the drum must be circular. Conversely, any deviation of this ratio from the maximum value signifies a non-circular shape. Moreover, the ratio of the first two tones cannot exceed $k_{1,1}^{2}/k_{0,1}^{2}$. This conclusively affirms the claim in Yau’s Problem 77 [48] for the case of an elastic drum.
While, for $\widetilde{R}\leq R$, we have the following conclusion.
Theorem 1.2. Let $\Omega\subset\mathbb{R}^{N}$ ($N\geq 2$) be a given bounded domain whose boundary satisfies an interior sphere condition. If $\widetilde{R}\leq R$, the ratio of the first two Robin eigenvalues of the Laplacian on $\Omega$ satisfies
$\frac{\mu_{2}}{\mu_{1}}\leq\frac{k_{\nu+1,1}^{2}-k_{\nu,1}^{2}}{\widetilde{R}^{2}\mu_{1}}+1$
for all $\sigma>0$, where $k_{\nu+l,m}$ denotes the $m$th positive zero of the function $kJ_{\nu+l+1}(k)-(\sigma+l)J_{\nu+l}(k)$. Furthermore, equality holds if and only if $\Omega=B_{R_{*}}$ with $R_{*}$ coming from Theorem 1.1.
If $\widetilde{R}<R$, one sees that
$\frac{\mu_{2}}{\mu_{1}}\Big|_{B_{R_{*}}}=\frac{k_{\nu+1,1}^{2}}{k_{\nu,1}^{2}}<\frac{k_{\nu+1,1}^{2}-k_{\nu,1}^{2}}{\widetilde{R}^{2}\mu_{1}}+1.$
At this point, our conclusion is slightly weaker than the conjecture proposed by Payne, Schaefer and Henrot. By Proposition 2.3 (in Section 2) and the right-continuity of $\mu(t)$, we have
$\lim_{\sigma\rightarrow+\infty}\mu\left(u_{1,pM}\right)=\lim_{\sigma\rightarrow+\infty}\mu\left(u_{1,m}\right)=|\Omega|,$
where $|\Omega|=\text{meas}\left(\Omega\right)$. If $\Omega$ is not a ball, in view of Lemma 3.2 (in Section 2), there exists $\sigma_{*}>0$ such that
$\left|B_{R}\right|\leq\mu\left(u_{1,pM}\right)$
for all $\sigma\geq\sigma_{*}$. This implies that $\widetilde{R}\geq R$. Therefore, for sufficiently large $\sigma$, the case $\widetilde{R}<R$ will not occur, which further implies that the above PPW type conjecture is valid for sufficiently large $\sigma$.
In fact, if $\Omega$ is not a ball, we have $\lambda_{2}(\Omega)/\lambda_{1}(\Omega)<\lambda_{2}\left(B_{R_{*}}\right)/\lambda_{1}\left(B_{R_{*}}\right)$ for the Dirichlet case, and hence, by the continuity, there exists $\sigma^{*}$ such that for any $\sigma>\sigma^{*}$ we have $\mu_{2}(\Omega)/\mu_{1}(\Omega)<\mu_{2}\left(B_{R_{*}}\right)/\mu_{1}\left(B_{R_{*}}\right)$ for the Robin case. This approach even not needs the interior sphere condition (Lipschitz is enough)  [^1] It is our pleasure to thank Professor Vladimir Bobkov for bringing this to our attention..
@@
Our results also provide a nuanced answer to Kac’s famous question [31], “Can one hear the shape of a drum?” While the general answer is negative, our theorems show that for an elastically supported drum, one can definitively “hear” its circularity from the first two tones.
@@
Theorem 1.3. The function $\sigma\mapsto k_{\nu+1,1}^{2}/k_{\nu,1}^{2}$ is continuously and strictly decreasing on $(0,+\infty)$, and its range spans from $\infty$ (as $\sigma\to 0^{+}$) down to $j_{\nu+1,1}^{2}/j_{\nu,1}^{2}$ (as $\sigma\to+\infty$).
@@
Figure 1 provides a schematic illustration of the ratio $k_{\nu+1,1}^{2}/k_{\nu,1}^{2}$ as a function of $\sigma$. A crucial implication of Theorem 1.3 for $N=2$ is the existence of a unique critical value $\overline{\sigma}>0$ such that $k_{\nu+1,1}^{2}/k_{\nu,1}^{2}\left(\overline{\sigma}\right)=3$. Furthermore, the ratio exceeds $3$ for $\sigma<\overline{\sigma}$ and falls below $3$ for $\sigma>\overline{\sigma}$. Consequently, when $\Omega$ is a ball, the inequality $\mu_{2}/\mu_{1}\leq 3$ holds if and only if $\sigma\geq\overline{\sigma}$. This result precisely identifies the optimal parameter range, thereby refining the earlier estimate in [39].
@@
where $\gamma=\sqrt{\lambda_{1}}/\alpha$. Chiti [13, 14, 15] proved, for a class of equations more general than the Laplacian, that there exists $s_{1}\in\left(0,\left|B_{1/\gamma}\right|\right)$ such that the decreasing rearrangements $\widetilde{u}_{1}^{*}(s)$ and $\widetilde{z}^{*}(s)$ (defined in Section 3) satisfy
$\left\{\begin{array}[]{ll}\widetilde{z}^{*}(s)\geq\widetilde{u}_{1}^{*}(s)&\text{for}\,\,\,\left[0,s_{1}\right],\\ \widetilde{z}^{*}(s)\leq\widetilde{u}_{1}^{*}(s)&\text{for}\,\,\,\left[s_{1},\left|B_{1/\gamma}\right|\right].\end{array}\right$.
This result was later extended by Ashbaugh and Benguria [7, Theorem A.2] to a broader class of second-order elliptic equations.
@@
The principal difficulty in deriving this inequality for the Robin problem, and a central challenge in applying symmetric rearrangement methods, stems fundamentally from the behavior of the eigenfunction on the boundary. Unlike the Dirichlet case, $u_{1}$ is generally non-constant on $\partial\Omega$. Consequently, the level set $\left\{x\in\Omega:u_{1}>t\right\}:=U_{t}$ may include portions of $\partial\Omega$ where $u_{1}>t$, which fundamentally alters its perimeter and geometric properties. This complication prevents a direct application of the techniques used for the fixed membrane problem [7, 13, 43], as noted in [2]. That’s why our conclusions depend on the maximum value of the eigenfunction on the boundary.
To overcome this challenge, we adapt an elegant idea from Alvino et al. [2, 3], which involves decomposing the boundary $\partial U_{t}$ of a level set into its internal part and its intersection with $\partial\Omega$. For $t\geq u_{1,pM}$, we may exclude the external part of $\partial U_{t}$. This yields the required isoperimetric-type inequality on a suitable subinterval, which is sufficient for our purposes. Building on this, we then derive a new Chiti-type comparison theorem. All these developments are presented in Section 3. Furthermore, leveraging the Faber-Krahn inequality for the Robin problem (established by Bossel [10] for $N=2$ and Daners [18] for $N\geq 2$), we also prove a key domain comparison result in this section, which serves as another essential prerequisite for the proof of Theorem 1.1.
The proof of Theorem 1.1 is presented in Section 4. Owing to the fundamental differences between our new Chiti-type comparison result and the classical version in [7, Theorem A.2], the structure of our proof diverges significantly from the original approach. In particular, establishing the necessity of the equality condition requires more complex and refined arguments. Section 5 presents the proof of Theorem 1.2. Given the structural similarities to Theorem 1.1, we focus primarily on elucidating the essential differences. Finally, the proof of Theorem 1.3 is completed in the last section.
@@ 2 Strict monotonicity of $w(r)$ and $B(r)$
where the first equality follows from the Bessel function identity $J_{\nu+1}(x)+J_{\nu+3}(x)=2(\nu+2)J_{\nu+2}(x)/x$, and the final inequality holds because $\beta<j_{\nu+1,1}$ ensures the integrand remains positive. ∎
@@
We conclude this section by analyzing the asymptotic behavior of the first eigenfunction with respect to the boundary parameter $\sigma$.
@@ 3 Chiti type comparison results
$\int_{\Omega}u^{2}(x)\,\text{d}x=\int_{0}^{|\Omega|}\left(u^{*}\right)^{2}(s)\,\text{d}s=\int_{\Omega^{*}}\left(u^{\star}\right)^{2}(x)\,\text{d}x,$ (3.1)
@@
Recall that $R=\alpha/\sqrt{\mu_{1}}$. Since $\alpha^{2}$ is the first eigenvalue of the Robin problem on the unit ball $B$, a standard scaling argument shows that $\mu_{1}$ is the first eigenvalue of the corresponding problem on the ball $B_{R}$
@@
where $c$ is a positive constant. Given that $\alpha=k_{\nu,1}$, we have $z(R)>0$, which contrasts with the Dirichlet case where the eigenfunction vanishes on the boundary.
@@
By the Faber-Krahn inequality for the Robin Laplacian, which was established by Bossel [10] for $N=2$ and Daners [18] for $N\geq 2$, we have
@@
It follows that $\mu_{1}\left(B_{R}\right)=\mu_{1}\left(\Omega^{*}\right)$. Thus one has that $B_{R}=\Omega^{*}$ which is a contradiction. Hence, $B_{R}\subseteq\Omega^{*}$ or $\left|B_{R}\right|\leq\left|\Omega^{*}\right|=|\Omega|$. Furthermore, if $\left|B_{R}\right|=|\Omega|$, then $B_{R}=\Omega^{*}$, and thus
@@
By Lemma 3.2, $\left|B_{R}\right|=|\Omega|$ if and only if $\Omega=B_{R}$. In this case, the conclusion of Theorem 1.1 holds trivially (by letting $R_{*}=R$) for all $\sigma>0$. We may therefore assume throughout the remainder of the paper that $\left|B_{R}\right|<|\Omega|$. Moreover, in the remaining part of this section, we always assume that $\widetilde{R}\geq R$.
@@
Lemma 3.3. Assume $\widetilde{R}\geq R$. For all $\sigma>0$, if $z^{*}(0)<u_{1}^{*}(0)$, then $z^{*}(s)<u_{1}^{*}(s)$ for all $s\in\left[0,\left|B_{R}\right|\right)$.
@@
Since $\widetilde{R}\geq R$, we derive that
@@
for all $\sigma>0$. By Lemma 3.1 and the definition of $v$, we have
@@
for almost every $s\in\left[0,s_{1}\right]$. Combining this with inequality (3.9), we conclude that the same inequality holds for almost every $s\in\left[0,\left|B_{R}\right|\right]$.
@@
$\int_{0}^{|\Omega|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s.$
@@
$\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s<\int_{0}^{\left|B_{R}\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s\leq\int_{0}^{\left|\Omega\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s,$ (3.10)
@@
Lemma 3.4. When $\widetilde{R}\geq R$ and $z^{*}(0)=u_{1}^{*}(0)$, the following holds: either $z^{*}(s)\geq u_{1}^{*}(s)$ for all $s\in\left[0,\left|B_{R}\right|\right]$, or there exists some $s_{1}\in\left(0,\left|B_{R}\right|\right)$ such that $z^{*}(s)\geq u_{1}^{*}(s)$ for $s\in\left[0,s_{1}\right]$ and $z^{*}(s)\leq u_{1}^{*}(s)$ for $s\in\left(s_{1},\left|B_{R}\right|\right]$.
@@
$\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s<\int_{0}^{\left|B_{R}\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s\leq\int_{0}^{\left|\Omega\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s,$
@@
Theorem 3.1. Let (3.5) hold and $\widetilde{R}\geq R$. For all $\sigma>0$, the functions $u_{1}^{*}(s)$ and $z^{*}(s)$ satisfy either
@@
(I) if $f=f(|x|)$ is nonnegative and decreasing (or increasing) as a function of $r=|x|$ for $x\in\Omega$, then $f^{\star}(r)\leq f(r)$ (or $f_{\star}(r)\geq f(r)$) for $r$ between 0 and the radius of $\Omega^{*}$, with strict inequality for $r\in\left(\min_{x\not\in}|x|,R_{*}\right)$ if $f$ is strictly monotone there, where $R_{*}$ denote the radius of $\Omega^{*}$;
@@
$f^{\star}(r)=f(r)\,\,\,\text{for}\,\,\,r\in\left[0,R_{*}\right].$
@@
$f_{\star}(r)=f(r)\,\,\,\text{for}\,\,\,r\in\left[0,R_{*}\right].$
@@
By Theorem 3.1, for all $\sigma>0$, the functions $u_{1}^{\star}(r)$ and $z(r)$ satisfy either
@@
$\left.-\int_{R}^{R_{*}}f(r)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right).$
@@
$\leq\,$ $NC_{N}\left(\int_{0}^{R}f(R)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{R_{*}}f(R)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $NC_{N}f(R)\left(\int_{0}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{R_{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$